total no. of voids in 0.5 mole of a compound forming hexagonal closed packed structure are

Grade:12th pass

1 Answers

Vignesh vijayakumar
6 years ago
0.5 mol contains (6.022×10^23)/2 particles=3.011×10^23 particlesNo. Of octahedral voids=no. Of particles=3.011×10^23-------( 1)No. Of tetrahedral voids =2×no. Of particles=6.022×10^23 -----------(2)Hence Total voids = (1)+(2)=(3.011+6.022)×10^23=9.033×10^23